OpenStudy (anonymous):

I need help with I think it's PMT the purchase cost is 250.000 and the interest rate is 6% the monthly year is 30 how do I find out what the monthly payments are? what the unpaid balance of the loan at the end of five years and the unpaid balance at the end of the 10 years help me please

OpenStudy (anonymous):

Can someone help me please

OpenStudy (anonymous):

So far I have PMT=250,000 (0.6) stuck after that

OpenStudy (anonymous):

hmmm.... basic rate where t is in months:\[C(t) = Co(1+ rate)^t \]\[C(t) = 250,000(1+0.06)^t \]\[C(t) = 250,000(1.06)^t\] 5 years:\[C(5) =250,000(1.06)^5 = 334,556\]334,556 - 250,000 = 84,556.40 10 years:\[C(10) =250,000(1.06)^10 = 447,712\]447,712 - 250,00 = 197,712 Is that right?

OpenStudy (anonymous):

So where did the one come from I don't hve that in each year

OpenStudy (anonymous):

it's from deriving the formula algebraically

OpenStudy (yuki):

pms?